# rowhighlight linewidth=1pt, linecolor=black, innertopmargin=4pt, innerbottommargin=4pt, innerleftmargin=4pt, innerrightmargin=4pt, backgroundcolor=lightgray!15

### v1

@@ Appendix B More Results on Synthetic
We further evaluate KronSAE on several variant block-diagonal covariance matrices (Figure 9). In each case, the decoder-weight covariance $C_{\mathrm{dec}}=W_{\mathrm{dec}}W_{\mathrm{dec}}^{\top}$ of KronSAE more faithfully reproduces the ground truth groupings than the TopK SAE. Notably, on the third covariance pattern (where some blocks are very small) TopK’s learned correlations nearly vanish, whereas KronSAE still uncovers the correct block structure. For the first covariance matrix, KronSAE yields sharply elevated correlations in the regions corresponding to the true blocks, in line with our design goal of head-wise compositionality.
Table 3 quantifies these observations via the RV coefficient and permutation tests. Even after optimally matching TopK’s atoms to a dense AE reference, TopK SAE attains only weak correlation alignment ($RV\approx 0.05-0.08$) with non-significant or marginal p-values. In contrast, KronSAE configurations achieve $RV$ values between 0.11 and 0.35 (all $p<0.001$), representing a 3–6x improvement in correlation recovery. These results confirm that our compositional encoder not only accelerates training but also robustly captures the intended hierarchical feature interactions across diverse covariance regimes.
Figure 9: Comparision of how different KronSAE try to reveal hidden structure of defined covariance matrix. The KronSAE models recover the underlying block-structured correlations more faithfully than the TopK baseline, with the finer head/composition split $(h=4,m=2)$ capturing smaller feature groups more accurately.
Covariance matrix SAE $RV$ p-value
1 TopK $0.046$ $0.1324$
TopK Matched $0.051$ $0.0102$
KronSAE $h=2,m=4,n=32$ $0.200$ $0.0002$
KronSAE $h=4,m=2,n=32$ $0.150$ $0.0002$
2 TopK $0.033$ $0.2212$
TopK Matched $0.080$ $0.0002$
KronSAE $h=2,m=4,n=32$ $0.221$ $0.0002$
KronSAE $h=4,m=2,n=32$ $0.319$ $0.0002$
3 TopK $0.035$ $0.2490$
TopK Matched $0.043$ $0.0002$
KronSAE $h=2,m=4,n=32$ $0.111$ $0.0002$
KronSAE $h=4,m=2,n=32$ $0.212$ $0.0002$
4 TopK $0.034$ $0.4613$
TopK Matched $0.043$ $0.0002$
KronSAE $h=2,m=4,n=32$ $0.346$ $0.0002$
KronSAE $h=4,m=2,n=32$ $0.334$ $0.0002$
Table 3: RV coefficient and permutation p-values for correlation recovery on four synthetic covariance patterns. KronSAE outperforms both the standard and atom-matched TopK SAE by a large margin, achieving statistically significant alignment ($p<10^{-3}$) across all cases.
Appendix C mAND as a Logical Operator
@@
For KronSAE, we replace the original $\operatorname{AND_{AIL}}$  [Lowe et al., 2021] with a more restrictive approximation, $\operatorname{mAND}$. Since our objective is to drive each atom toward a distinct, monosemantic feature, we found that tightening the logical conjunction encourages sharper feature separation. Moreover, by using the geometric mean ($\sqrt{p\,q}$) rather than a simple product or minimum, $\operatorname{mAND}$ preserves activation magnitudes. A visual comparison of $\operatorname{mAND}$ and $\operatorname{AND_{AIL}}$ appears in Figure 10.
@@
Figure 10: Comparison of the smooth $\operatorname{mAND}$ operator against the $\operatorname{AND_{AIL}}$  [Lowe et al., 2021].
@@ Appendix D Feature Analysis Methodology
We analyze learned features using an established pipeline Bills et al. [2023], Paulo et al. [2024] consisting of three stages: (1) statistical property collection, (2) automatic activation pattern interpretation, and (3) interpretation evaluation. The following subsections detail our implementation.
@@ D.1 Data collection
multitoken $\text{multitoken ratio}=\frac{1}{b}\sum_{i=0}^{b}\frac{\text{number of activations in sequence $i$}}{\text{total tokens in sequence $i$}},$ (7)
@@ D.2 Feature interpretations
For each feature, we generate interpretations by sampling 16 random activation examples above the median activation quantile and presenting them to Qwen3 14B [Yang et al., 2025] (AWQ-quantized with reasoning enabled). The model produces concise descriptions of the activation patterns. Empirical observations suggest reasoning mode improves interpretation quality, though we lack quantitative measurements. This aligns with findings in [Paulo et al., 2024], which compared standard one-sentence responses with Chain-of-Thought outputs, making model reasoning an interesting direction for future research.
@@ D.3 Evaluation pipeline
We evaluate interpretations using balanced sets of up to 64 positive (activation quantile > 0.5) and 64 negative examples, employing the same model without reasoning to reduce computational costs. When insufficient examples exist, we maintain class balance by equalizing positive and negative counts. The evaluation uses modified system prompts from [Paulo et al., 2024], with added emphasis on returning Python lists matching the input example count exactly. We discard entire batches if responses are unparseable or contain fewer labels than the number of provided examples.
@@
score=12(correctly predicted positivestotal positives+correctly predicted negativestotal $\text{score}=\frac{1}{2}\left(\frac{\text{correctly predicted positives}}{\text{total positives}}+\frac{\text{correctly predicted negatives}}{\text{total negatives}}\right).$ (8)
@@ Appendix E Additional Feature Analysis Results
As noted in Section 5.3, latent embeddings primarily exhibit clustering within their originating groups (head, base, extension). With the support of observations reported in Sections 4.2 and 5.3, we find that models with more heads achieve better reconstruction while producing more diverse basis vectors. This suggests that fine-grained architectures yield more expressive representations, although they may also exhibit undesired challenging behavior like feature splitting [Bricken et al., 2023] or absorption [Chanin et al., 2024].
@@ Appendix F KronSAE in Terms of Tensor Diagram
The proposed encoder architecture can be visualized as a tensor diagram (Figure 16). Notably, this formulation draws a connection to quantum mechanics, where |$\ket{\mathbf{f}}$ represents the (unnormalized) state of two disentangled qubits described by |$\ket{\boldsymbol{p}}$ and |$\ket{\boldsymbol{q}}$.
If we were to sum the outputs of the encoder’s heads instead of concatenating them, |$\ket{\mathbf{f}}$ would correspond to a separable quantum state. This scenario can be expressed via the Schmidt decomposition:
|$\ket{\mathbf{f}}=\sum_{h}\ket{\boldsymbol{p}_{h}}\otimes_{K}\ket{\boldsymbol{q}_{h}},$
@@
Figure 16: For a single head, the KronSAE encoder architecture separates the input $\boldsymbol{x}$ into two distinct components, $\boldsymbol{p}$ and $\boldsymbol{q}$, via matrix multiplications with $W_{enc}^{p}$ and $W_{enc}^{q}$ accordingly. These components are then combined via the outer product $\boldsymbol{p}\otimes\boldsymbol{q}$, resulting in a matrix representation. To produce the final output vector $\mathbf{f}$, this matrix is flattened into a single vector using a multi-index mapping.
@@ Appendix G Analysis of Compositional Structure
Latents presented in Table 6 emonstrates more consistent AND-like behavior than Heads 3 and 136, closely matching the interaction pattern shown in Figure 2, supporting our theoretical framework.

### v3

@@ Appendix B More Results on Synthetic
In this section we present the motivation behind our matching algorithm and additional results.
B.1 Feature Matching as Quadratic Problem
Suppose that we have two sets of feature embeddings from different models represented as matrices $X,Y\in\mathbb{R}^{F\times d}$, where $F$ is the number of features and $d$ is the dimensionality of feature embeddings. Our task is to find the optimal assignment between features from $Y$ to features from $X$ so that this assignment would satisfy some considerations.
The standard approach is to solve linear assignment problem - find a permutation matrix $\Pi$ subject to minimizing the $\text{trace}(C^{T}\Pi)$, where $C$ is a cost matrix defined as pairwise distance between features $C_{i,j}=d(X_{i},Y_{j})$. Standard algorithm for solving it in context of sparse dictionary learning is a Hungarian algorithm (paulo2025sparseautoencoderstraineddata; balagansky2025mechanistic; fel2025archetypal).
This linear problem only considers pairwise information between features while ignoring the global dependencies between features within $X$ and $Y$ sets separately, e.g. clusters should map to clusters, and the linear problem does not internalize this information. Our observation of correlations between features naturally requires to search for assignment that would take this information into account: we seek for a permutation $\Pi$ that would give best global alignment, measured as the Frobenius norm of $X^{T}\Pi Y$. So the objective becomes:
$\max_{\Pi}\|X^{T}\Pi Y\|^{2}_{F}=\max_{\Pi}\text{trace}(\Pi^{T}XX^{T}\Pi YY^{T}),$ (8)
where global feature structure is explicitly encoded in the matrices $XX^{T}$ and $YY^{T}$.
An efficient algorithm to solve the quadratic assignment problem is Fast Approximate Quadratic Programming (FAQ) method (Vogelstein2015FastAQ) that initially was designed for graphs matching: given the adjacency matrices $A$ and $B$, it minimizes the $\text{trace}(A\Pi B^{T}\Pi^{T})$, where $\Pi$ are relaxed from permutation matrices to the set of doubly stochastic matrices (Birkhoff polytope). In our case we define $A=XX^{T}$ and $B=YY^{T}$, and since we do not want to minimize the cost but rather maximize the similarity, we solve for the reversed objective:
$\max_{\Pi}(XX^{T}\Pi(YY)^{T}\Pi^{T})=\max_{\Pi}\text{trace}(\Pi^{T}XX^{T}\Pi YY^{T}),$ (9)
which is the same as equation 8, and this formulation preserves global dependencies because the contribution of assigning $Y_{j}$ to $X_{i}$ depends on all other assignments through the cross-terms in the quadratic form. Listing LABEL:listing:assignment shows the implementation of this matching procedure.
Listing 1: Implementation of FAQ algorithm for quadratic feature assignment problem.
1def feature_matching(A, B, max_iter):
2 F, d = A.shape[0]
3 G_A, G_B = A @ A.T, B @ B.T
4 P = np.ones((F, F)) / F # Barycenter initialization
5
6 # Frank-Wolfe iterations
7 for _ in range(max_iter):
8 grad = 2 * G_A @ P @ G_B # Gradient for maximization
9 r, c = linear_sum_assignment(-grad) # Solve LAP
10 Q = np.zeros_like(P)
11 Q[r, c] = 1
12
13 # Compute optimal step size
14 D = Q - P
15 b, a = np.trace(grad.T @ D), np.trace(G_A @ D @ G_B @ D.T)
16
17 if abs(a) < 1e-12:
18 alpha = 1.0 if b > 0 else 0.0
19 elif a < 0: # Concave case
20 alpha = np.clip(-b/(2*a), 0, 1)
21 else: # Convex case
22 alpha = 1.0 if b > 0 else 0.0
23
24 P_new = P + alpha * D
25 P = P_new
26
27 # Project to permutation matrix
28 r, c = linear_sum_assignment(-P)
29 P_final = np.zeros((F, F))
30 P_final[r, c] = 1
31
32 return P_final, P_final @ B
B.2 Additional Results
As shown in Figure 13, KronSAE maintains correlation structure that is heavier than in TopK and is better aligned with ground truth covariance matrix.
Figure 13: Examples of patterns learned in autoencoder TopK and KronSAE variants after we apply the improved matching scheme. KronSAE have learned patterns that more closely resemble the underlying ground truth structure, and with increasing number of heads (more fine-grained structure) it approaches the TopK SAE.
Table 7 indicate that structure learned by KronSAE is more diverse as indicated by standard deviation of the corresponding values, and is also more aligned with the ground truth covariance according to higher RV scores and lower differences ($\Delta$) between properties of ground truth and learned matrix.
Model RV coeff. Effective rank Mean corr. Rank $\Delta$
AE 0.59 $\pm$ 0.17 63.3 $\pm$ 0.3 0.13 $\pm$ 0.04 47.9
TopK 0.12 $\pm$ 0.02 58.5 $\pm$ 1.2 0.08 $\pm$ 0.01 43.2
Kron $h=2,m=4,n=32$ 0.22 $\pm$ 0.07 53.8 $\pm$ 3.4 0.08 $\pm$ 0.01 38.5
Kron $h=4,m=2,n=32$ 0.3 $\pm$ 0.12 46.0 $\pm$ 7.8 0.10 $\pm$ 0.03 30.7
Kron $h=4,m=8,n=8$ 0.23 $\pm$ 0.08 48.8 $\pm$ 5.0 0.09 $\pm$ 0.02 33.5
Kron $h=8,m=2,n=16$ 0.17 $\pm$ 0.04 55.9 $\pm$ 1.7 0.07 $\pm$ 0.01 40.6
Table 7: Results with improved matching scheme, computed for 8 different covariance setups. Higher RV coefficients between ground truth and learned matrices and lower $\Delta$ between properties of these matrices indicate that KronSAE is more variable across different setups and is better aligned with ground truth correlation structures.
Together, these results additionally validate improved covariance reconstruction in KronSAE.
Appendix C mAND as a Logical Operator and KronSAE as Logical SAE
AND-like mechanism and hierarchy. Suppose we have $\mathbf{u}^{k}=P^{k}\mathbf{x}$ and $\mathbf{v}^{k}=Q^{k}\mathbf{x}$. KronSAE can be described in two equivalent ways:
  1. 1.
    Applying the $\operatorname{ReLU}$ to latents $\mathbf{u}$, $\mathbf{v}$, then applying the Kronecker product and square root.
  2. 2.
    Applying the $\operatorname{mAND}$ kernel that creates the matrix as in equation 3 and flattening it.
These approaches are complementary. To formally understand how they induce the AND-like mechanism and hierarchy, consider base and extension pre-latent vectors $\mathbf{p}=\operatorname{ReLU}(\mathbf{u})$ and $\mathbf{q}=\operatorname{ReLU}(\mathbf{v})$. Each $p_{i}$ is a $i$th base pre-latent activation, and $q_{j}$ is $j$th extension pre-latent activation. Kronecker product $\mathbf{p}\otimes\mathbf{q}$ creates the vector $(p_{1}*q_{1},…,p_{1}*q_{m},p_{2}*q_{1},…,p_{n}*q_{m})$ of the activations of post-latents. Then there is two situations:
  1. 1.
    Post-latent is active $\implies$ both pre-latents activations are positive.
  2. 2.
    Post-latent is inactive $\implies$ at least one of pre-latent activations is zero.
Hence post-latent is active only when both pre-latents are active. Fix some post-latent and its corresponding pre-latents, and suppose that $\mathcal{P},\mathcal{Q},\mathcal{F}$ are the sets of input vectors from the hidden state space (passed to SAE) on which base pre-latent, extension pre-latent and post-latent are active. Then it is true that $\mathcal{F}=\mathcal{P}\cap\mathcal{Q}$, meaning that pre-latents must be broader and polysemous to encode multiple semantics of emitted post-latents. We validate this behaviour qualitatively in section 5.3, although in the same section we describe that apparently other types of interactions also presented.
Matryoshka and Kron hierarchy. In contrast to TopK SAE, Matryoshka loss imposes different kind of hierarchy by dividing the dictionary into groups of $G_{1},…,G_{k}$ latents where each is of different level of granularity. Namely, first $G_{1}$ latents are the most broad and abstract, next $G_{2}$ latents add more fine-grained semantics, lowering the level of abstraction, and so on. This type of structure imposed by specific loss function - increasing the level of granularity must decrease the reconstruction error, and the lowest level of $G_{1}$ features must also maintain good reconstruction quality - and does not strictly demand some kind of conditional activation of features, while KronSAE imposes two-level AND-like hierarchy via architectural design (Kronecker product). As we have two different mechanisms of feature hierarchy (from encoding mechanism and loss function design) we can combine it, as shown in Section 4.1, 4.2 and Appendix A.4, to combine properties of both approaches.
@@ B.2 Additional Results
Visual intuition. We also compare our $\operatorname{mAND}$ to existing $\operatorname{AND_{AIL}}$  (Lowe2021). Since our objective is to drive each atom toward a distinct, monosemantic feature, we found that tightening the logical conjunction encourages sharper feature separation. Moreover, by using the geometric mean ($\sqrt{p\,q}$) rather than a simple product or minimum, $\operatorname{mAND}$ preserves activation magnitudes and prevents post-latent activation to be exploded when both $p,q$ are positive. A visual comparison of $\operatorname{mAND}$ and $\operatorname{AND_{AIL}}$ appears in Figure 14.
@@
Figure 14: Comparison of the smooth $\operatorname{mAND}$ operator against the $\operatorname{AND_{AIL}}$  (Lowe2021).
@@ Appendix D Feature Analysis Methodology
We analyze learned features using an established pipeline bills2023llmcanexplain; paulo2024automatically consisting of three stages: (1) statistical property collection, (2) automatic activation pattern interpretation, and (3) interpretation evaluation. The following subsections detail our implementation.
@@ D.1 Data collection
[…]
@@ D.2 Feature interpretations
For each feature, we generate interpretations by sampling 16 random activation examples above the median activation quantile and presenting them to Qwen3 14B (yang2025qwen3technicalreport) (AWQ-quantized with reasoning enabled). The model produces concise descriptions of the activation patterns. Empirical observations suggest reasoning mode improves interpretation quality, though we lack quantitative measurements. This aligns with findings in (paulo2024automatically), which compared standard one-sentence responses with Chain-of-Thought outputs, making model reasoning an interesting direction for future research.
@@ D.3 Evaluation pipeline
We evaluate interpretations using balanced sets of up to 64 positive (activation quantile > 0.5) and 64 negative examples, employing the same model without reasoning to reduce computational costs. When insufficient examples exist, we maintain class balance by equalizing positive and negative counts. The evaluation uses modified system prompts from (paulo2024automatically), with added emphasis on returning Python lists matching the input example count exactly. We discard entire batches if responses are unparseable or contain fewer labels than the number of provided examples.
@@
$\text{score}=\frac{1}{2}\left(\frac{\text{correctly predicted positives}}{\text{total positives}}+\frac{\text{correctly predicted negatives}}{\text{total negatives}}\right).$ (12)
@@ Appendix E Additional Feature Analysis Results
As noted in Section 5.3, latent embeddings primarily exhibit clustering within their originating groups (head, base, extension). With the support of observations reported in Sections 4.1 and 5.3, we find that models with more heads achieve better reconstruction while producing more diverse basis vectors. This suggests that fine-grained architectures yield more expressive representations, although they may also exhibit undesired challenging behavior like feature splitting (bricken2023towards) or absorption (chanin2024absorption).
@@
Interpretability across sparsity regimes.
We compare KronSAE and its Matryoshka variant with TopK, Matryoshka and Switch SAE baselines across different sparsity regimes for Gemma 2 2B and Qwen 2.5 1.5B. For each SAE we follow the same pipeline as for results in Section 5.3, but use only 18 million tokens for examples collection. For KronSAE and TopK we use first 4096 features, and for other models we sample 4096 features randomly. Scores for $F=2^{16}$ and $F=2^{15}$ are presented in Figures 20 and 21.
Figure 20: Interpretability scores for $F=2^{16}=65536$. KronSAE shows good consistency across different sparsity regimes, in some cases outperforming the TopK baseline.
Figure 21: Interpretability scores for $F=2^{15}=32768$. KronSAE shows good consistency across different sparsity regimes, in some cases outperforming the TopK baseline. The result is the same as for 65k dictionary size at the Figure 20, but with slightly lower scores.
Results show that TopK and KronSAE are very stable across both models and sparsity regimes, while for other models their architectural and training design significantly affect the interpretability.
We attribute those differences between Gemma and Qwen to different capacity of residual stream - Gemma has hidden state size of 2304, while Qwen has only 1536 dimensions (1.5x smaller). Hovewer, differences between training data and architectural choices can also be the cause.
Interpretability tradeoff.
As shown in Section 4.1 KronSAE introduces tradeoff between computational efficiency, explained variance and interpretability. We can force more features to be correlated in same head by increasing the $m$ and $n$, and this will improve computational efficiency, but at the cost of reconstruction performance; however, improving the explained variance (with small $m$ and fine-grained structure of groups) goes with the cost of slight reduction in the interpretability. These tradeoffs are expected and presented across variety of SAEs (karvonen2025saebench).
To evaluate how KronSAE behave under different $m,n$ we compute autointerpretability scores following the same setup as described in Appendix D for Qwen-2.5 1.5B and Gemma-2 2B models with sparsity $\ell_{0}\in\{16,32,64,128\}$. As Figure 22 shows, setup $m=1$ and $h=4096$ is less interpretable than $m=2$ and $m=4$ with the same number of heads despite having stronger reconstruction performance.
Figure 22: EV, absorption score and interpretability for $F=2^{16}$. For constant $h$, decreasing $m$ leads to lower interepretability scores and absorption scores, but increase EV - this is expected due to increasing polysemanticity of pre-latents and entanglement of the features. See Section 5.3 for a explanation of the retrieval mechanism.
@@ Appendix F KronSAE in Terms of Tensor Diagram
[…]
@@
Figure 23: For a single head, the KronSAE encoder architecture separates the input $\bm{x}$ into two distinct components, $\bm{p}$ and $\bm{q}$, via matrix multiplications $P^{k}$ and $Q^{k}$ accordingly, followed by $\operatorname{ReLU}$ activation. These components are then combined via the Kronecker product $\bm{p}\otimes\bm{q}$ and square root operation $\sqrt{\cdot}$
, resulting in an output vector $\mathbf{f}$.
@@ Appendix G Analysis of Compositional Structure
Latents presented in Table 10 emonstrates more consistent AND-like behavior than Heads 3 and 136, closely matching the interaction pattern shown in Figure 2.
@@
Appendix H KronSAE Simplified Implementation
1class KronSAE($n$.Module):
2 def __init__(self, config):
3 super().__init__()
4 self.config = config
5 _t = torch.nn.init.normal_(
6 torch.empty(
7 self.config.act_size,
8 self.config.h * (self.config.m + self.config.n)
9 )
10 ) / math.sqrt(self.config.dict_size * 2.0)
11 self.W_enc = $n$.Parameter(_t)
12 self.b_enc = $n$.Parameter(
13 torch.zeros(self.config.h * (self.config.m + self.config.n))
14 )
15 W_dec_v0 = einops.rearrange( # Initialize decoder weights
16 _t.t().clone(), "(h mn) d -> h mn d",
17 h=self.config.h, mn=self.config.m + self.config.n
18 )[:, :self.config.m]
19 W_dec_v1 = einops.rearrange(
20 _t.t().clone(), "(h mn) d -> h mn d",
21 h=self.config.h, mn=self.config.m + self.config.n
22 )[:, self.config.m:]
23 self.W_dec = $n$.Parameter(einops.rearrange(
24 W_dec_v0[..., None, :] + W_dec_v1[..., None, :, :],
25 "h m n d -> (h m n) d"
26 ))
27 self.W_dec.data[:] = (
28 self.W_dec.data / self.W_dec.data.norm(dim=-1, keepdim=True)
29 )
30 self.b_dec = $n$.Parameter(torch.zeros(self.config.act_size))
31
32 def encode(self, x: torch.Tensor) -> torch.Tensor:
33 B, D = x.shape
34 acts = F.relu(
35 x @ self.W_enc + self.b_enc
36 ).view(B, self.h, self.m + self.n)
37 all_scores = torch.sqrt(
38 acts[..., :self.config.m, None] * \
39 acts[..., self.config.m:, None, :] + 1e-5
40 ).view(B, -1)
41 scores, indices = all_scores.topk(
42 self.config.k, dim=-1, sorted=False
43 )
44 acts_topk = torch.zeros(
45 (B, self.config.dict_size)
46 ).scatter(-1, indices, scores)
47 return acts_topk
48
49 def forward(self, x):
50 acts_topk = self.encode(x)
51 x_rec = acts_topk @ self.W_dec + self.b_dec
52 output = self.get_loss_dict(x, x_rec)
53 return output
54
55 def get_loss_dict(self, x, x_rec):
56 loss = (x_rec - x.pow(2).mean()
57 pt_l2 = (x_rec - x).pow(2).sum(-1).squeeze()
58 var = (x - x.mean(0)).pow(2).sum(-1).squeeze()
59 ev = (1 - pt_l2 / var).mean()
60 return loss, ev
Appendix I Usage of Large Language Models
We have used LLMs as the main tool for conducting the interpretability experiments, as described in section D, and as the instrument for language polishing and word choice.